OpenStudy (anonymous):

A mercury lamp emits radiation with a wavelength of 4.36 X 10^-7 m. (a) What is the wavelength of this radiation in centimeters? (b) In what region of the electromagnetic spectrum is this radiation? (c) Calculate the frequency of this radiation.

OpenStudy (anonymous):

for part a, it is just a simple conversion of m to cm. so multiply by 100 b. look it up online c. can't help

OpenStudy (anonymous):

so for part a) i got 6.9x10^6 m; how to convert to cm?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aaronq):

how did you get that for part a)? it's wrong. it's just a conversion; 1 m = 100 cm b) Look it up online for c) use: \(\lambda*\nu=c\)

OpenStudy (anonymous):

actually it was right. i looked it up online in the teachers edition book @aaronq

OpenStudy (aaronq):

"A mercury lamp emits radiation with a wavelength of 4.36 X 10^-7 m. (a) What is the wavelength of this radiation in centimeters?" you said "6.9x10^6 m" it can't be a different number. 1 m = 100 cm \(4.36 *10^{-7} m=4.36 *10^{-7} *\dfrac{100~cm}{1~m}=4.36 *10^{-5}~cm\)